OpenStudy (anonymous):

Please help! I am so lost on this one. A. Line m contains the points, A(–2, 6) and B(4, 8), while line n contains the C(8, 12) and D(x, 24). Given m and n are perpendicular lines, solve for the value of x. Given m and n are parallel lines, solve for the value of x. In your final answer, include all formulas and calculations necessary to solve for x.

OpenStudy (s4sensitiveandshy):

find the slope of both line slope formula \[m= \frac{y_2-y_1}{x_2-x_1}\] \[(x_1,y_1)(x_2,y_2)\] A(–2, 6) and B(4, 8) x_1=-2 x_2=4 y_16 y_2=8 replace x's and y's with the given coordinaes

OpenStudy (s4sensitiveandshy):

coordinates

OpenStudy (s4sensitiveandshy):

Perpendicular lines: the slope of perpendicular lines should be *negative reciprocal* if x/y is the slope of the original equation then the perpendicular slope would be *-y/x* and parallel lines must have the same slope.

OpenStudy (s4sensitiveandshy):

so.. what is the slope of line m ?

OpenStudy (anonymous):

I am so confused , So I would do 8 squared -16/4 squared --2?

OpenStudy (s4sensitiveandshy):

that is not squared y_2 means 2nd y-coordinate \[(x_1,y_1) =(1st ~x-coordinate , 1st~y-coordinate )\] \[(x_2 ,y_2)= (2nd ~c-coordinate , 2nd~y-coordinate )\]

OpenStudy (s4sensitiveandshy):

 so replace x_1 with -2 and y_1 with 6

OpenStudy (s4sensitiveandshy):

do you mean \[\frac{ 8-6 }{ 4--2 }\] ?

OpenStudy (anonymous):

WOW I get it now ... I was so lost I confused myself.

OpenStudy (anonymous):

m1 = (8 - 6)/(4 - (-2)) m1 = 2/6 m1 = 1/3

OpenStudy (s4sensitiveandshy):

nice so what would be the *perpendicular slope * ??

OpenStudy (anonymous):

-1?

OpenStudy (s4sensitiveandshy):

hmm no perpendicular slopes are negative reciprocals for example 2/3 is the perpendicular slope of -3/2

OpenStudy (s4sensitiveandshy):

flip the fraction and change the sign. that is it. :)

OpenStudy (anonymous):

So it would be 6/2 see this is what confused me

OpenStudy (anonymous):

Sorry to be so slow

OpenStudy (s4sensitiveandshy):

you need perpendicular slope of *1/3* flip the fraction, change sign what would you get ?

OpenStudy (anonymous):

x=4?

OpenStudy (s4sensitiveandshy):

how did you get that ? :)

OpenStudy (anonymous):

Sorry lost the page, nothing but issues today. It was 1/3*12/8-8=-1 -4-x =4=x

OpenStudy (s4sensitiveandshy):

i have to go. but ... use the same formula to find slope of line n \[\frac{ 24-12 }{ x-8 }=m\] * Given m and n are perpendicular lines, solve for the value of x* replace m with the perpendicular slope to find x value for the 2nd part * Given m and n are parallel lines, solve for the value of x* replace m with the parallel slope and then solve for x.

OpenStudy (anonymous):

Thank you for the help!!!

OpenStudy (s4sensitiveandshy):

that is not correct 8-8 =0 \[\frac{ 12 }{ 0 }=undefined \] you can't divide by 0.

OpenStudy (s4sensitiveandshy):

if 1/3 is the original equation then what would be the perpendicular slope can you answer this question,please ?

OpenStudy (anonymous):

2/6

OpenStudy (s4sensitiveandshy):

m1 = (8 - 6)/(4 - (-2)) m1 = 2/6 m1 = 1/3 slope is 1/3 right ? perpendicular slopes are * negative reciprocal* so flip the fraction and change the sign 1/3 perpendicular slope would be -3/1 which is same as -3

OpenStudy (s4sensitiveandshy):

\[\frac{ 24-12 }{ x-8 }=m\] * Given m and n are perpendicular lines, solve for the value of x* replace m with the perpendicular slope to find x value so replace m with -3 to find x value for parallel lines replace m with 1/3 (because parallel lines must have the same slope )

OpenStudy (s4sensitiveandshy):

hope that make sense now. practice!